Part 2 of 4 Required information NOTE: This is a multi-part question. Once an answer is submitted, you will be unable to retua) The temperature of the steam at the turbine exit

b) Thermal efficiency of the cycle
c) the mass flow rate of the steam
Answer #1


at P, = 10kPa and aio ni = 19118 KIlig Na 0.00101 m3 kg wp-i = 4, (P2 - P.) 5000 OM Misen) Pump 10kPa 0.00101 (0000 - 10) - 0

0.75 = 347901 – no 347901 - 2461-1 no=2715-6 kilig at Po= 10 kPa q ho = 2715-6 Koles - To = 115c a) temperature of steam. To